OpenStudy (anonymous):

(2x²+16x+30)/(5x²+13-6) Simplify

OpenStudy (ganpat):

can u factorize them ??

OpenStudy (anonymous):

no im not really good at that

OpenStudy (ganpat):

no probs..

OpenStudy (ganpat):

2x²+16x+30 divide throughout by 2.. so, x2 + 8x + 15 the factors r 5 & 3... as 5 * 3= 15 and, 5 + 3 = 8 x2 + 5x + 3x + 15 = 0 x (x +5) + 3 (x +5) =0 (x + 5) (x + 3).. can u try the same for other equation ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thank u

OpenStudy (ganpat):

5x²+13-6 5x2 + 15x -2x - 6 = 0 5x (x +3) -2 (x +3) = 0 (5x -2 ) (x +3) = 0.. Now dividing both the equation, as given i question.. (x + 5) (x + 3) / (5x -2 ) (x +3 = (x +5) / (5x -2) .. is the answer..

OpenStudy (kaiz122):

@Ganpat , since u have divided it by 2, you need to multiply it again by 2 the factor for the numerator would be 2(x+5)(x+3)

OpenStudy (ganpat):

oh yes.. my bad.. sorry.. it shd be multiplied back.. SO the answer = 2(x +5) / (5x -2)